The quantity (3^5/4)(9^3/2) can be written in the form 3^x. What is the value of x?
Answer by rothauserc(4718)

(3^5/4)(9^3/2) = 3^(5/4) * (3^3) = 3^(17/4)
therefore x = 17/4